Consider the following information from the market for lemonade: Price Quantity Demanded Quantity Supplied $1 500 cups 150 cups $2 200 cups 310 cups

(a) As the price changes from $1 to $2, what is the value of price elasticity of demand?
1
Expert's answer
2021-10-13T10:18:16-0400

I shall use midpoint approach for calculating elasticity

a)

Initial quantity"=Q_o=500"

Final quantity"=Q_1=200"

Average quantity"=Q=\frac{(500+200)}{2}=350"

Initial price"=P_o=\$1"

Final price"=P_o=\$2"

Average price"=Q=\frac{(1+2)}{2}=\$1.50"


Price elasticity of demand"=\frac{[\frac{(Q_1-Q_o)}{Q}]}{[\frac{(P_1-P_o)}{P}]}=\frac{[\frac{(200-500)}{350}]}{[\frac{(2-1)}{1.50}]}=-1.28571"